OpenStudy (anonymous):

Express y as a function of x. lny=7x+lnC

jimthompson5910 (jim_thompson5910):

the goal is to isolate y, so you need to undo that ln (natual log)

jimthompson5910 (jim_thompson5910):

what undoes the natural log? ie what is the inverse of a natural log?

OpenStudy (anonymous):

The number e?

jimthompson5910 (jim_thompson5910):

yep, e^x is the inverse

jimthompson5910 (jim_thompson5910):

lny=7x+lnC e^( lny )=e^( 7x+lnC ) y = e^(7x + lnC) y = e^(7x) * e^(lnC) y = e^(7x) * C y = C * e^(7x)

jimthompson5910 (jim_thompson5910):

hopefully that makes sense